A solution of propylene glycol (antifreeze) in the car's radiator has a freezing point of -20°C. What is the molality of the solution?

1
Expert's answer
2022-03-31T15:09:02-0400

ΔTf = Kf * m

Kf - molal freezing point depression constant or cryoscopic constant

m - molality of the solution


ΔTf = Tf0 - Tfs

Tf0  represents the freezing point of the pure solvent

Tfs  represents the freezing point of the solution

ΔT = T(H2O) - T(C3H8O2)

T(C3H8O2) = -20°C = 253.15K

ΔT = 273K - 253.15K=19.85K


K(H2O) = 1.86 K*kg/m


m = ΔT / K(H2O) = 19.85 K/ 1.86 K*kg/m = 10.67 kg/m